# mddefDefinition

### v2

@@ A.1 Derivation of Boltzmann Estimator
We note that a similar estimator is obtained in Thaler & Zavadlav (2021) by differentiating through a reweighting scheme arising from thermodynamic perturbation theory. We have presented an alternative derivation that does not require reweighting.
@@ A.3 StABlE Training Algorithm
2:Pre-trained Neural Network Potential $U_{\theta}$
@@ A.4 Observables
where $\delta$ is the Dirac-Delta function. Although the observable need not be differentiable with respect to $r$ for our Boltzmann learning framework, we compute a differentiable version of $h(r)$ via Gaussian smearing as in Wang et al. (2023a) to facilitate comparison with differentiable simulation methods.
@@ A.5 Stability Criteria
Adapted from Fu et al. (2022), the maximum bond length deviation metric captures unphysical bond stretching or collapse in small flexible molecules. According to this criterion, a simulation becomes unstable at time $T$ if,
@@
where $\mathcal{B}$ is the set of all bonds, $i,j$ are the two endpoint atoms of the bond, and $b_{i,j}$ is the equilibrium bond length computed from the reference simulation. Following Fu et al. (2022), we set $\Delta=0.5A$ for final stability evaluation. However, we adopt a more conservative value of $\Delta=0.25A$ during training in order to detect and correct instability earlier. We use this criterion for the MD17 and MD22 datasets.
@@
The minimum intermolecular distance metric is appropriate at train-time to detect instability early before it cascades to the rest of the system. However, it is too sensitive to use for evaluation, as realistic simulation can still be achieved for some time after the occurrence of a highly localized instability. Therefore, following Fu et al. (2022), we adopt an instability metric based on the radial distribution function, defined as,
@@ A.6 Architecture and Training Details
We provide details on the model architectures and training procedures used in this work. MD simulations and MLFFs are written in the PyTorch framework and are built upon the MDsim Fu et al. (2022) and Atomic Simulation Environment Larsen et al. (2017) packages. All training is performed on a single NVIDIA A100 GPU.
@@
SchNet Schütt et al. (2018) E(3)-invariant 0.12M 5.0 -
NequIP Batzner et al. (2022) E(3)-equivariant 0.12M 5.0 1
GemNet-T Gasteiger et al. (2021) SE(3)-equivariant 1.89M 5.0 -
@@
For energy and forces pre-training, we follow the protocols in Fu et al. (2022). In order to isolate the effect of observable-based learning, we begin StABlE Training only after pre-training has fully converged (that is, when $\mathcal{L}_{\text{QM}}$ has reached a plateau). This means that any improvements in stability or accuracy as a result of StABlE Training can be attributed to the learning signal from the reference observable, as opposed to the regularization from the QM energy and forces data.
@@ A.7 Simulation Details
We provide MD simulation details in Supplementary Table 4. During training, all systems are simulated with a Nose-Hoover thermostat. During evaluation, either a Nose-Hoover or Langevin thermostat is used based on whichever one yields better stability. Thermostat parameters are chosen to be consistent with prior literature Chmiela et al. (2017; 2022); Fu et al. (2022): for Nose-Hoover simulations, the temperature coupling constant is set to 20 fs, and for Langevin simulations, the friction coefficient is set to 0.1 ps-1.
@@ A.8 Evaluation Details
Following Fu et al. (2022), we choose stability thresholds such that a realistic, high-fidelity simulation at the chosen temperature would virtually never cross the threshold. This means that if a simulation does cross the threshold, this is indicative of catastrophic failure. Thresholds are set more conservatively during training in order to facilitate early detection of potential collapse. Supplementary Figure 7 shows the distribution of values of the stability criterion over high-fidelity reference simulations for the three systems considered in this work, along with thresholds chosen to denote instability for training and evaluation.
@@ A.9 Temperature-Reweighting of Observables
We can then compute a reweighted Monte Carlo estimate of the observable at $T_{2}$ as follows Thaler & Zavadlav (2021).
@@
The statistical error of the reweighted Monte Carlo estimate is captured by the effective sample size, $N_{eff}\approx e^{-\sum_{i=1}^{N}w_{i}ln(w_{i})}$ Carmichael & Shell (2012). A small effective sample size indicates that a few samples with high weights dominate the average; this occurs for large differences between $T_{1}$ and $T_{2}$. To select lower and upper temperatures at which to perform the temperature generalization experiment, we set a minimum $N_{eff}=1000$, leading us to choose 350K and 700K (Figure 8).
@@ A.11 Analysis of Energy and Forces Errors
Finally, we note that the observed energy MAE increase on aspirin for our chosen combination of learning rate ($\alpha=0.001$) and loss coefficient ($\lambda=10$) is from 0.87 to 1.4 $\mathrm{kcal\,mol}^{-1}$, while DFT error for energies on MD17 can be as high as 2.3 $\mathrm{kcal\,mol}^{-1}$ Faber et al. (2017). Thus, some of the error in the MLFF predictions could be attributable to inaccuracies in the underlying DFT data.
@@ A.12 Effect of Simulation Timestep on Stability
We perform 100 ps simulations using various timesteps for the aspirin and Ac-Ala3-NHMe tetrapeptide systems, using a SchNet and NequIP potential respectively (Supplementary Figure 11).
Figure 11: Effect of timestep on simulation stability. Reducing the timestep does not completely eliminate instability, and can sometimes worsen stability.
We observe that instability is not completely eliminated as the timestep is reduced. For the tetrapeptide system, stability consistently decreases as the timestep is reduced. Similar behavior has been observed in neural network based solvers for ordinary differential equations Krishnapriyan et al. (2023). For the aspirin system, stability improves as the timestep is reduced, but does so very slowly (simulations are not completely stable even with a timestep of 0.05 fs , which is 10 times lower than the original timestep).
A.13 Velocity Autocorrelation Function of Aspirin

### v3

@@ A.1 Derivation of Boltzmann Estimator
We note that a similar estimator is obtained in (Thaler & Zavadlav, 2021) by differentiating through a reweighting scheme arising from thermodynamic perturbation theory. We have presented an alternative derivation that does not require reweighting.
@@ A.3 StABlE Training Algorithm
2:Pre-trained Machine Learning Force Field $U_{\theta}$
@@ A.4 Observables
where $\delta$ is the Dirac-Delta function. Although the observable need not be differentiable with respect to $r$ for our Boltzmann learning framework, we compute a differentiable version of $h(r)$ via Gaussian smearing as in (Wang et al., 2023a) to facilitate comparison with differentiable simulation methods.
@@ A.5 Stability Criteria
Adapted from (Fu et al., 2022), the maximum bond length deviation metric captures unphysical bond stretching or collapse in small flexible molecules. According to this criterion, a simulation becomes unstable at time $T$ if,
@@
where $\mathcal{B}$ is the set of all bonds, $i,j$ are the two endpoint atoms of the bond, and $b_{i,j}$ is the equilibrium bond length computed from the reference simulation. Following (Fu et al., 2022), we set $\Delta=0.5A$ for final stability evaluation. However, we adopt a more conservative value of $\Delta=0.25A$ during training in order to detect and correct instability earlier. We use this criterion for the MD17 and MD22 datasets.
@@
The minimum intermolecular distance metric is appropriate at train-time to detect local instability early before it cascades to the rest of the system. However, it is too sensitive to use for evaluation, as realistic simulation can still be achieved for some time after the occurrence of a highly localized instability. Therefore, following (Fu et al., 2022), we adopt an instability metric based on the radial distribution function, defined as,
@@ A.6 Architecture and Training Details
We provide details on the model architectures and training procedures used in this work. MD simulations and MLFFs are written in the PyTorch framework and are built upon the MDsim (Fu et al., 2022) and Atomic Simulation Environment (Larsen et al., 2017) packages. All training is performed on a single NVIDIA A100 GPU.
@@
SchNet (Schütt et al., 2018) E(3)-invariant 0.12M 5.0 -
NequIP (Batzner et al., 2022) E(3)-equivariant 0.12M 5.0 1
GemNet-T (Gasteiger et al., 2021) SE(3)-equivariant 1.89M 5.0 -
@@
For energy and forces pre-training, we follow the protocols in (Fu et al., 2022). In order to isolate the effect of observable-based learning, we begin StABlE Training only after pre-training has fully converged (that is, when $\mathcal{L}_{\text{QM}}$ has reached a plateau). This means that any improvements in stability or accuracy as a result of StABlE Training can be attributed to the learning signal from the reference observable, as opposed to the regularization from the QM energy and forces data.
@@
General Guidelines for Choosing Hyperparameters.
$t$ should be chosen large enough that the deviation of ensemble averages computed within the window from ground truth values are primarily attributable to systematic error/physical instability rather than sampling error. If t is chosen too large, the frequency of gradient updates reduces, slowing down learning. Generally, a frequency of 1 picosecond should be sufficient for structural observables of small molecular systems, and may need to be larger (10-100 ps) for larger-scale or coarse-grained systems. The number of replicas $R$ should be chosen so as to maximize MLFF inference throughput (samples/second) while remaining within GPU memory. Since we perform simulations in parallel by vectorizing over the batch dimension, we see steady improvements in throughput until GPU memory saturates, at which point performance plateaus or degrades. The minibatch size $B$ should also be chosen as large as possible to minimize variance in the Boltzmann Estimator, while remaining within GPU memory limits.
@@ A.7 Simulation Details
We provide MD simulation details in Supplementary Table 4. During training, all systems are simulated with a Nose-Hoover thermostat. During evaluation, either a Nose-Hoover or Langevin thermostat is used based on whichever one yields better stability. Thermostat parameters are chosen to be consistent with prior literature (Chmiela et al., 2017; 2022; Fu et al., 2022): for Nose-Hoover simulations, the temperature coupling constant is set to 20 fs, and for Langevin simulations, the friction coefficient is set to 0.1 ps-1.
@@ A.8 Evaluation Details
Following (Fu et al., 2022), we choose stability thresholds such that a realistic, high-fidelity simulation at the chosen temperature would virtually never cross the threshold. This means that if a simulation does cross the threshold, this is indicative of catastrophic failure. Thresholds are set more conservatively during training in order to facilitate early detection of potential collapse. Supplementary Figure 7 shows the distribution of values of the stability criterion over high-fidelity reference simulations for the three systems considered in this work, along with thresholds chosen to denote instability for training and evaluation. As a rough guideline for new systems, we suggest setting the threshold at 4 standard deviations beyond mean fluctuations for training, and 5 standard deviations for evaluation.
@@ A.9 Temperature-Reweighting of Observables
We can then compute a reweighted Monte Carlo estimate of the observable at $T_{2}$ as follows (Thaler & Zavadlav, 2021).
@@
The statistical error of the reweighted Monte Carlo estimate is captured by the effective sample size, $N_{eff}\approx e^{-\sum_{i=1}^{N}w_{i}ln(w_{i})}$ (Carmichael & Shell, 2012). A small effective sample size indicates that a few samples with high weights dominate the average; this occurs for large differences between $T_{1}$ and $T_{2}$. To select lower and upper temperatures at which to perform the temperature generalization experiment, we set a minimum $N_{eff}=1000$, leading us to choose 350K and 700K (Figure 8).
@@ A.11 Analysis of Energy and Forces Errors
Finally, we note that the observed energy MAE increase on aspirin for our chosen combination of learning rate ($\alpha=0.001$) and loss coefficient ($\lambda=10$) is from 0.87 to 1.4 $\mathrm{kcal\,mol}^{-1}$, while DFT error for energies on MD17 can be as high as 2.3 $\mathrm{kcal\,mol}^{-1}$ (Faber et al., 2017). Thus, some of the error in the MLFF predictions could be attributable to inaccuracies in the underlying DFT data.
As rough guidelines for new systems, if Force MAE is prioritized, then the learning rate should be smaller and the QM loss coefficient should be set higher. If stability improvements are prioritized over Force MAE, such as in cases where the reference energy/force data is known to be unreliable, the opposite is true.
@@ A.12 Effect of Simulation Timestep on Stability
We perform 100 ps simulations with 32 replicates, using various timesteps for the aspirin and Ac-Ala3-NHMe tetrapeptide systems, using a SchNet and NequIP potential respectively (Supplementary Figure 11).
Figure 11: Effect of reducing timestep on simulation stability. Reducing the timestep does not completely eliminate instability, and can sometimes worsen stability.
We observe that instability is not completely eliminated as the timestep is reduced. For the tetrapeptide system, stability consistently decreases as the timestep is reduced. Similar behavior has been observed in neural network based solvers for ordinary differential equations (Krishnapriyan et al., 2023). For the aspirin system, stability improves as the timestep is reduced, but does so very slowly (simulations are not completely stable even with a timestep of 0.05 fs , which is 10 times lower than the original timestep).
We also investigate the effect of increasing beyond the original timestep of 0.5 fs on simulation stability for Aspirin and Ac-Ala3-NHMe. We again perform 100 ps simulations with 32 independent replicates, now with timesteps of 1, 2, 5, and 10 fs. We observe that StABlE Training yields stability improvements at larger timesteps up to 2 fs, but after this point, neither the pretrained nor StABlE-trained potential are able to simulate stably for an appreciable amount of time (Supplementary Figure 12). We emphasize that we cutoff the simulations at 100 ps, so the aspirin simulation with the StABlE-trained model using a timestep of 1 fs would likely simulate stably for considerably longer if not cut off.
Figure 12: Effect of increasing timestep on simulation stability. StABlE Training improves simulation stability for timesteps up to 2 fs, after which stability rapidly deteriorates.
A.13 Comparison of Boltzmann Estimator to Alternative Differentiation Strategies
We compare our Boltzmann estimator with two alternative differentiation strategies, namely direct backpropagation through the unrolled MD simulation, and the adjoint method described in (Chen et al., 2019). As in (Wang et al., 2023a), we consider a system with 32 particles governed by a Lennard-Jones potential acting on the pairwise particle distances. We initialize the simulations with a prior potential capturing only the repulsive term of the potential, and seek to learn a correction term, parameterized by a multi-layer perceptron with 5 hidden layers of size 128, so as to reproduce the behavior of the full potential. We utilize supervision from the ground truth radial distribution function. We measure the loss gradient norms, memory footprint, and runtime of all approaches as a function of the simulation length, showing results in Figure 13. As expected, direct backpropagation quickly runs out of memory because it needs to store intermediate network activations after every forward pass. The adjoint method eliminates this memory requirement by performing a backwards ODE solve to calculate the loss gradients. However, as reported in (Wang et al., 2023a), the adjoint dynamics are highly unstable over long rollouts and lead to exploding gradient norms. Meanwhile, our Boltzmann estimator achieves roughly constant gradient norms as the simulation length increases due to the decoupling of the gradient computation from the dynamics, and also has a favorably low memory and compute footprint.
Figure 13: Comparison of Boltzmann Estimator to direct backpropagation and adjoint method on a toy Lennard-Jones system . The Boltzmann Estimator achieves stable gradient norms and favorable memory and runtime footprints as the simulation length is increased. Direct backpropagation is memory prohibitive, and the adjoint method suffers from unstable dynamics, eventually causing gradient norms to explode.
A.14 Velocity Autocorrelation Function of Aspirin